OpenStudy (anonymous):

A building has a ramp to its front doors to accommodate the handicapped. If the distance from the building to the end of the ramp is 17 feet and the height from the ground to the front doors is 7 feet, how long is the ramp? (Round to the nearest tenth.)

OpenStudy (anonymous):

Use Pythagorean Therom a^2 +b^2 = c^2

OpenStudy (anonymous):

haha I was gonna say that

OpenStudy (anonymous):

7^2+17^2=c^2 c^2=338

OpenStudy (anonymous):

\[\sqrt{c^2} = \sqrt{338}\]

OpenStudy (anonymous):

C = 18.384776310850235634421953414726

OpenStudy (anonymous):

OK i got it could you help with a couple problems please ? Just explain it and show me how you got the answer if you do you got a fan .